# Are LLMs Good Annotators for Discourse-level Event Relation Extraction?

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated proficiency in a wide array of natural language processing tasks. However, its effectiveness over discourse-level event relation extraction (ERE) tasks remains unexplored. In this paper, we assess the effectiveness of LLMs in addressing discourse-level ERE tasks characterized by lengthy documents and intricate relations encompassing coreference, temporal, causal, and subevent types. Evaluation is conducted using an commercial model, GPT-3.5, and an open-source model, LLaMA-2. Our study reveals a notable underperformance of LLMs compared to the baseline established through supervised learning. Although Supervised Fine-Tuning (SFT) can improve LLMs performance, it does not scale well compared to the smaller supervised baseline model. Our quantitative and qualitative analysis shows that LLMs have several weaknesses when applied for extracting event relations, including a tendency to fabricate event mentions, and failures to capture transitivity rules among relations, detect long distance relations, or comprehend contexts with dense event mentions.
@@
Kangda Wei, Aayush Gautam, Ruihong Huang Department of Computer Science and Engineering, Texas A&M University {kangda, aayushgautam}@tamu.edu, huangrh@cse.tamu.edu
@@ 1 Introduction
  • Transitivity and hallucinations: There is a discernible tendency for LLMs to violate the rules of transitivity among event relation predictions. Furthermore, LLMs display inconsistencies in adhering to the provided prompts. Both suggest a potential lack of reasoning abilities and accurate understanding on the assigned task.
@@ 2 Related Works
Event relation extraction (ERE) has been one of the fundamental challenges for natural language processing Chaturvedi et al. (2017); Rashkin et al. (2018); Zhang et al. (2020). As understanding relations between events is crucial for understanding human languages Levelt (1989); Miller and Johnson-Laird (1976) and beneficial for various applications Khashabi et al. (2019); Zhang et al. (2020); Choubey and Huang (2018), many approaches have been developed for performing ERE tasks Liu et al. (2014); Hashimoto et al. (2014); Ning et al. (2017), and many high performing approaches are based on supervised learning Dligach et al. (2017); Aldawsari and Finlayson (2019); Liu et al. (2020); Lu and Ng (2021). However, few-shot LLMs have not been sufficiently explored for ERE tasks.
@@
Many recent works have studied how to make LLMs perform better through applying various prompting techniques, including role-prompting Zhang et al. (2023); Buren (2023), re-sampling Holtzman et al. (2020), one-shot or few-shot prompting au2 et al. (2021); Shyr et al. (2023), and question decomposition Wei et al. (2023c) etc. Other novel and advanced techniques include Chain of Thought prompting Wei et al. (2023a), least-to-most prompting Zhou et al. (2023), learning from multiple prompts Wei et al. (2023b), and retrieval augmentation Lazaridou et al. (2022); Jiang et al. (2023). We refer to these techniques as guidelines when designing prompts in this work.
@@ 3.1 Data
For this study, we use the MAVEN-ERE dataset created by Wang et al. (2022), which includes annotations of four types of event relations: coreference, temporal, causal, and subevent. MAVEN-ERE consists of 4480 English Wikipedia documents, containing $103,193$ event coreference chains, $1,216,217$ temporal relations, $57,992$ causal relations, and $15,841$ subevent relations. MAVEN-ERE is challenging as the documents contain comprehensive relation types, event relations at the discourse level and have denser relations among events comparing to other datasets. For example, MAVEN-ERE has an average of 272 temporal relation links per document comparing to 49 temporal relation links per document for MATRES Ning et al. (2018); TimeBank-Dense Cassidy et al. (2014) mainly focus on sentence-level relations; and TDDiscourse Naik et al. (2019) only consider temporal relations.
@@ 3.3 Model
To get an idea of how the newest GPT model performs, we also tested gpt-4-1106-preview model on a subset of the validation set instead of the whole test set becuase API calls to GPT-4 models are 30 times more expensive than GPT-3.5 models [^5] Performance reported in Appendix E.
@@ 4 Results
Iterative MUC B3 CEAFe BLANC
@@
Do LLMs still underperform supervised-learning method when fine-tuned with training data? We answer this question in Figure 3 by varying the amount of training data for the supervised baseline method and meanwhile fine-tuning LLaMA-2999Here, LLaMA-2 is fine-tuned in the pairwise format as described in Section 3.2.4 to match with the baseline model Wang et al. (2022) which treats the ERE tasks as pairwise classification tasks. Detailed hyperparameters for fine-tuning LLaMA-2 can be found in Appendix A. with the same amount of training data.
SFT certainly improves the performance of LLaMA-2. However, it still underperforms the supervised baseline method for all the relation types as the number of used training documents increases. LLaMA-2 typically requires twice more training data to reach the same overall performance as the smaller supervised baseline model. LLaMA-2 only outperforms the supervised baseline method when the available training data is very limited, which benefits from its zero-shot capability emergent from large-scale pre-training. It also deserves mentioning that fine-tuning LLMs is much more expensive than fine-tuning smaller language models like RoBERTa. For example, fine-tuning using 200 training documents for 3 epochs requires approximately 72 hours for LLaMA-2 7B, while only about 1 hour is needed for the roberta-base baseline model.
@@ 5 Discussion
According to Wang et al. (2022), 88.8% temporal relations and 23.9% causal relations can be inferred with transitivity rules in the MAVEN-ERE dataset. Failure to follow the transitivity rules detriments GPT’s performance. Moreover, Table 3 highlights that false negatives and false positives emerge as the predominant error types, indicating that GPT-3.5 faces challenges in accurately discerning the presence or absence of relationships. Notably, in the context of temporal relationships, a noteworthy increase in the F1 score is observed when incorporating all transitivity-inferred relations. This implies that GPT-3.5 indeed falls short of capturing a comprehensive array of transitive relations.

### v3

@@ Abstract
Large Language Models (LLMs) have demonstrated proficiency in a wide array of natural language processing tasks. However, its effectiveness over discourse-level event relation extraction (ERE) tasks remains unexplored. In this paper, we assess the effectiveness of LLMs in addressing discourse-level ERE tasks characterized by lengthy documents and intricate relations encompassing coreference, temporal, causal, and subevent types. Evaluation is conducted using an commercial model, GPT-3.5, and an open-source model, LLaMA-2. Our study reveals a notable underperformance of LLMs compared to the baseline established through supervised learning. Although Supervised Fine-Tuning (SFT) can improve LLMs performance, it does not scale well compared to the smaller supervised baseline model. Our quantitative and qualitative analysis shows that LLMs have several weaknesses when applied for extracting event relations, including a tendency to fabricate event mentions, and failures to capture transitivity rules among relations, detect long distance relations, or comprehend contexts with dense event mentions. Code available at: https://github.com/WeiKangda/LLM-ERE.git.
@@
Kangda Wei, Aayush Gautam, Ruihong Huang Department of Computer Science and Engineering Texas A&M University, College Station, TX {kangda, aayushgautam, huangrh}@tamu.edu
@@ 1 Introduction
  • Transitivity and hallucinations: There is a discernible tendency for LLMs to violate the rules of transitivity among event relation predictions. Furthermore, LLMs display inconsistencies in adhering to the provided prompts. Both suggest a potential lack of reasoning abilities and inaccurate understanding on the assigned task.
@@ 2 Related Works
Event relation extraction (ERE) has been one of the fundamental challenges for natural language processing Chaturvedi et al. (2017); Rashkin et al. (2018); Zhang et al. (2020). As understanding relations between events is crucial for understanding human languages Levelt (1989); Miller and Johnson-Laird (1976) and beneficial for various applications Khashabi et al. (2019); Zhang et al. (2020); Choubey and Huang (2018), many approaches have been developed for performing ERE tasks Liu et al. (2014); Hashimoto et al. (2014); Ning et al. (2017), and many high performing approaches are based on supervised learning Dligach et al. (2017); Aldawsari and Finlayson (2019); Liu et al. (2020); Lu and Ng (2021).
@@
Many recent works have studied how to make LLMs perform better through applying various prompting techniques, including role-prompting Zhang et al. (2023); Buren (2023), re-sampling Holtzman et al. (2020), one-shot or few-shot prompting au2 et al. (2021); Shyr et al. (2023), and question decomposition Wei et al. (2023c). Other novel and advanced techniques include Chain of Thought prompting Wei et al. (2023a), least-to-most prompting Zhou et al. (2023), learning from multiple prompts Wei et al. (2023b), and retrieval augmentation Lazaridou et al. (2022); Jiang et al. (2023). We refer to these techniques as guidelines when designing prompts in this work.
@@ 3.1 Data
For this study, we use the MAVEN-ERE dataset created by Wang et al. (2022), which includes annotations of four types of event relations: coreference, temporal, causal, and subevent. MAVEN-ERE consists of 4480 English Wikipedia documents, containing $103,193$ event coreference chains, $1,216,217$ temporal relations, $57,992$ causal relations, and $15,841$ subevent relations.
MAVEN-ERE is challenging as the documents contain comprehensive relation types, event relations at the discourse level and have denser relations among events comparing to other datasets. For example, MAVEN-ERE has an average of 272 temporal relation links per document comparing to 49 temporal relation links per document for MATRES Ning et al. (2018); TimeBank-Dense Cassidy et al. (2014) mainly focus on sentence-level relations; and TDDiscourse Naik et al. (2019) only consider temporal relations.
@@ 3.3 Model
To get an idea of how the newest GPT model performs, we also tested gpt-4-1106-preview model on a subset of the validation set instead of the whole test set becuase API calls to GPT-4 models are 30 times more expensive than GPT-3.5 models. The model performance was reported in Appendix E.
@@ 4 Results
Iterative MUC Vilain et al. (1995) B3 Bagga and Baldwin (1998) CEAFe Luo (2005) BLANC Recasens and Hovy (2011)
@@
Will LLMs perform better when we fine-tune them with training data? How will LLMs compare to the smaller baseline model when fine-tuned using the same amount of data? We answer this question by varying the amount of training data for the supervised baseline method and meanwhile fine-tuning LLaMA-2with the same amount of training data. In this experiment, LLaMA-2 is fine-tuned in the pairwise format as described in Section 3.2.4 to match with the baseline model Wang et al. (2022) which treats the ERE tasks as pairwise classification tasks. Detailed hyperparameters for fine-tuning LLaMA-2 can be found in Appendix A.
Figure 3 shows the result comparisons between LLaMA-2 and the smaller baseline model. First, SFT certainly improves the performance of LLaMA-2. However, it still underperforms the smaller baseline method for all the relation types as the number of used training documents increases. LLaMA-2 typically requires twice more training data to reach the same overall performance as the smaller baseline model. LLaMA-2 only outperforms the smaller baseline method when the available training data is very limited, as LLaMA-2 benefits from its zero-shot capability emergent from large-scale pre-training. It also deserves mentioning that fine-tuning LLMs is much more expensive than fine-tuning the smaller baseline language model RoBERTa. For example, fine-tuning using 200 training documents for 3 epochs requires approximately 72 hours for LLaMA-2 7B, while only about one hour is needed for the roberta-base baseline model.
@@ 5 Discussion
According to Wang et al. (2022), 88.8% temporal relations and 23.9% causal relations can be inferred with transitivity rules in the MAVEN-ERE dataset. Failure to follow the transitivity rules detriments GPT’s performance. Moreover, Table 3 highlights that false negatives and false positives emerge as the predominant error types, indicating that GPT-3.5 faces challenges in accurately discerning the presence or absence of relationships. Notably, in the context of temporal relationships, a noteworthy increase in the F1 score is observed when incorporating all transitivity-inferred relations. This implies that GPT-3.5, indeed falls short of capturing a comprehensive array of transitive relations.
@@ 6 Conclusion
Acknowledgements
We would like to thank the anonymous reviewers for their valuable feedback and input. We gratefully acknowledge support from National Science Foundation via the award IIS-1942918. Portions of this research were conducted with the advanced computing resources provided by Texas A&M High-Performance Research Computing.